OpenStudy (anonymous):

if log of 2 to the base 10=.301 log of 25 to the base 8=? show how to solve

OpenStudy (dumbcow):

\[\log_{10} 2 = .301 \] \[\log_{8} 25 = \frac{\log 25}{\log 8} = \frac{1.398}{.903} = 1.548 \] Basically use change of base formula \[\log_{b} x = \frac{\log x}{\log b} \]

OpenStudy (anonymous):

hah! you can't do that?

OpenStudy (anonymous):

\[\log_{8}25=2\log_{10}5/3\log_{10}2=2(1-\log_{10}2)/3\log_{10}2=2*(1-.301)/(3*.301)=1.548 \]

OpenStudy (dumbcow):

very nice :)